OpenStudy (anonymous):

How can I write an equation that is perpendicular to y=3x+1 that passes through (0,4)?

OpenStudy (confluxepic):

This site can help you. https://www.desmos.com/calculator If two lines are perpendicular then the product of their slopes is -1.

OpenStudy (anonymous):

How do I find the slope of the line?

OpenStudy (anonymous):

without using a website

OpenStudy (confluxepic):

This site could help you in finding slopes. http://coolmath.com/algebra/08-lines/06-finding-slope-line-given-two-points-01.htm

OpenStudy (confluxepic):

\[Slope=\frac{ Rise }{ Run }\]

OpenStudy (confluxepic):

m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

really

OpenStudy (anonymous):

You have to use the equation y = mx + b m is the slope number You insert your 0,4 (x,y) cordinates for the y and the x of the equation and then you solve

OpenStudy (anonymous):

@allyr I think it would be y= mx + 4. is that right?

OpenStudy (confluxepic):

You can get two points by finding the x-intercept and y-intercept.

OpenStudy (anonymous):

i really don't understand

OpenStudy (anonymous):

I would say yes but I am only certainly familiar with the equation of a line that is parallel to another and passes through a point so I don't know for sure about it being perpendicular

TheSmartOne (thesmartone):

http://prntscr.com/4xvf42 lol

OpenStudy (anonymous):

hilarious